OpenStudy (anonymous):

check my answers? 1. what equation represents the following word sentence? the quotient of a number b and 0.3 equals negative 10 i think it's b/0.3 = -10 2. which equation is equivalent to the following equation? -3/4x + 1/8 = -3/8 a. -3/4x = -3/8 - 1/8 b. -3/4x = -3/8 + 1/8 c. x + 1/8 = -3/8 * (-4/3) d. x + 1/8 = -3/8 * (-3/4) i think it's b

OpenStudy (anonymous):

3. a student was solving the following equation: -96 = -6(x - 15) -96 = -6x - 90 -96 + 90 = -6x - 90 + 90 -6 = -6x -6/-6 = -6x/-6x 1 = x what should they do to correct the error that they made? a. first add 6 to both sides b. first subtract x from both sides c. distribute the -6 to get 6x - 90 d. distribute the -6 to get -6x + 90 i think it's c

OpenStudy (anonymous):

write an expression in simplest form that represents the perimeter of the polygon. the measure of the sides are 3x, x + 10, 2x + 2, and 5x. would the simplified expression be 11x + 12?

OpenStudy (anonymous):

Number one looks good. For number two i would say A because if you take a negative to the other side of an equal sign it will be positive. Well for number three i think it's D because when you open -6(x-15) that is equal to -6x+90.

OpenStudy (anonymous):

Hope i helped. : )

OpenStudy (anonymous):

thanks, i still need help with that last one though

OpenStudy (anonymous):

The last one looks fine to me.

OpenStudy (anonymous):

okay thanks again

OpenStudy (anonymous):

No problem. Any time.